# The separating Noether number of small groups

### v2

@@ 7. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
Case 2.a: $q_{yy}(w_{2})=0$ and $p^{-}_{yyy}(w_{2})\neq 0$. Then $w_{2}$ is a non-zero scalar multiple of $[1,0]^{T}$ or $[0,1]^{T}$, say $w_{2}$ is a non-zero scalar multiple of $[1,0]^{T}$. If $w_{1}$ is a scalar multiple of $[0,1]^{T}$, then none of $q_{xy}$, $q^{-}_{xy}$, $q_{xy}q^{-}_{xy}$, $q_{xy}^{2}$, $p^{-}_{yyy}$ vanishes on $v=(w_{1},w_{2})$, and these relative invariants on $V$ of degree at most $4$ exhaust all the five non-trivial weights, so we are done. Otherwise the first coordinate of $w_{1}$ is non-zero, and thus $p_{xxy}$, $p_{xyy}$, $p^{-}_{xxy}$, $p^{-}_{xyy}$ are relative invariants on $V$ of degree $3$ not vanishing on $v$ (just like $p^{-}_{yyy}$). So we have exhausted all the five non-trivial weights, and we are done.
Case 2.b: $p^{-}_{yyy}(w_{2})=0$ and $q_{yy}(w_{2})\neq 0$. Then $w_{2}$ is a non-zero scalar multiple of $[1,1]^{T}$, $[1,\omega]^{T}$, or $[1,\omega^{2}]^{T}$. Moreover, $w_{1}$ is not a scalar multiple of $w_{2}$ (since otherwise its stabilizer would contain the stabilizer of $w_{2}$). Then $q_{yy}$, $q^{-}_{xy}$, $q_{yy}^{2}$, $q^{-}_{xy}q_{yy}$, $p^{-}_{xyy}$ are non-zero on $v=(w_{1},w_{2})$, and these relative invariants on $W_{1}\oplus W_{2}$ of degree at most $4$ exhaust all the five non-trivial weights.
Case 3: $V=W_{1}\oplus W_{3}$. The group $G$ has the automorphism $\alpha:a\mapsto a$, $b\mapsto b$, $c\mapsto c^{2}$. We have $\psi_{3}=\psi_{2}\circ\alpha$, so $\alpha$ interchanges $W_{2}$ and $W_{3}$, and permutes the $1$-dimensional representations. So Case 3 follows from Case 2 by Lemma 3.11. ∎
@@
8. Some groups $G=\mathrm{C}_{n}\rtimes\mathrm{C}_{k}$ with $\beta_{\mathrm{sep}}(G)<\beta^{K}(G)$
8.1. The group $\mathrm{C}_{4}\rtimes\mathrm{C}_{4}$
In this section
$G=\mathrm{C}_{4}\rtimes\mathrm{C}_{4}=\langle a,b\mid a^{4}=1=b^{4},\quad bab^{-1}=a^{3}\rangle.$
Assume that $K$ has an element $\mathrm{i}$ of multiplicative order $4$. Then $G$ has two irreducible $2$-dimensional representations (up to isomorphism), namely
$\psi_{1}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix},\qquad\psi_{2}:a\mapsto\begin{bmatrix}\mathrm{i}&0\\ 0&-\mathrm{i}\\ \end{bmatrix},\quad b\mapsto\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix}.$
The other irreducible representations of $G$ are $1$-dimensional, and can be labeled by the elements of the subgroup $\widehat{G}:=\{\pm 1\}\times\{\pm 1,\pm\mathrm{i}\}$ of $K^{\times}\times K^{\times}$ as follows: identify $\chi=(\chi_{1},\chi_{2})\in\widehat{G}$ with the representation $\chi:G\to K^{\times}$ given by
$\chi:a\mapsto\chi_{1},\quad b\mapsto\chi_{2}.$
Denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$ for $j=1,2$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$. Set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$. Note that none of $\psi_{1}$ or $\psi_{2}$ is faithful: $\ker(\psi_{1})=\langle b^{2}\rangle$, $G/\ker(\psi_{1})\cong\mathrm{D}_{8}$, and $\psi_{1}$ is the lift to $G$ of the only $2$-dimensional representation of $\mathrm{D}_{8}$. On the other hand, $\ker(\psi_{2})=\langle a^{2}b^{2}\rangle$, $G/\ker(\psi_{2})\cong\mathrm{Dic}_{8}$, and $\psi_{2}$ is the lift to $G$ of the only $2$-dimensional irreducible representation of the quaternion group. The coordinate functions on $W_{1}$ are denoted by $x_{1},x_{2}$, the coordinate functions on $W_{2}$ are $y_{1},y_{2}$, and the coordinate function on $U_{\chi}$ is $t_{\chi}$ for $\chi\in\widehat{G}$. It is well known and easy to see that
$K[x_{1},x_{2}]^{G}$ $=K[x_{1}x_{2},x_{1}^{4}+x_{2}^{4}],$
$K[y_{1},y_{2}]^{G}$ $=K[(y_{1}y_{2})^{2},\ y_{1}^{4}+y_{2}^{4},\ y_{1}y_{2}(y_{1}^{4}-y_{2}^{4})].$
Consider the following relative invariants (their index indicates their weight):
$r_{(-1,1)}:=x_{1}^{2}+x_{2}^{2},\quad r_{(-1,-1)}:=x_{1}^{2}-x_{2}^{2},\quad r_{(1,-1)}:=r_{(-1,1)}r_{(-1,-1)}=x_{1}^{4}-x_{2}^{4}$
$s_{(1,-1)}:=y_{1}y_{2},\quad s_{(-1,1)}:=y_{1}^{2}+y_{2}^{2},\quad s_{(-1,-1)}:=y_{1}^{2}-y_{2}^{2}.$
Using the above observations, one can prove by arguments similar to the proof of Proposition 6.1 the following:
Proposition 8.1.
For $i=1,2$ we have $\beta(G,W_{i}\oplus U)=6$.
Proposition 8.2.
We have $\beta(G,W_{1}\oplus W_{2})\leq 6$.
Proof.
We know that $\beta^{K}(G)\leq 7$ by [15, Proposition 3.1]. On the other hand, the element $a^{2}\in G$ acts on $W_{1}\oplus W_{2}$ as multiplication by the scalar $-1$, hence any homogeneous element of $K[W_{1}\oplus W_{2}]^{G}$ has even degree. It follows that $\beta(G,W_{1}\oplus W_{2})$ is an even number, less than or equal to $7$. Consequently, $\beta(G,W_{1}\oplus W_{2})\leq 6$. ∎
Proposition 8.3.
We have $\beta_{\mathrm{sep}}(W_{1}\oplus W_{2}\oplus U_{(1,\mathrm{i})})\leq 6$.
Proof.
Let $v=(w_{1},w_{2},u)\in V$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$ such that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}$ with $\deg(f)\leq 6$ (here $w_{i},\ w^{\prime}_{i}\in W_{i}$ and $u,u^{\prime}\in U$). We need to show that $v$ and $v^{\prime}$ have the same $G$-orbit. By Proposition 8.2, $(w_{1},w_{2})$ and $(w^{\prime}_{1},w^{\prime}_{2})$ are in the same $G$-orbit. Replacing $v$ by an appropriate element in its $G$-orbit, we may assume that $w_{1}=w^{\prime}_{1}$ and $w_{2}=w^{\prime}_{2}$. If $w_{1}=0$ or $w_{2}=0$, then we have $G\cdot v=G\cdot v^{\prime}$ by Proposition 8.1. From now on we assume that none of $w_{1}$ or $w_{2}$ is zero. Set
$r^{(1)}_{(1,-\mathrm{i})}:=x_{1}y_{2}+\mathrm{i}x_{2}y_{1},\quad r^{(2)}_{(1,-\mathrm{i})}:=x_{1}y_{1}^{3}-\mathrm{i}x_{2}y_{2}^{3},\quad r_{(1,\mathrm{i})}:=x_{1}y_{2}-\mathrm{i}x_{2}y_{1}.$
Then $r^{(1)}_{(1,-\mathrm{i})}$, $r^{(2)}_{(1,-\mathrm{i})}$, $t^{3}$ are relative invariants of weight $(1,-\mathrm{i})\in\widehat{G}$, whereas $r_{(1,\mathrm{i})}$, $t$ are relative invariants of weight $(1,\mathrm{i})$. Therefore we have the invariants
$f_{1}:=t^{4},\quad f_{2}:=r^{(1)}_{(1,-\mathrm{i})}t,\quad f_{3}:=r^{(2)}_{(1,-\mathrm{i})}t,\quad f_{4}:=r_{(1,\mathrm{i})}t^{3}.$
All of the above invariants have degree at most $5$. By assumption, $f_{j}(v)=f_{j}(v^{\prime})$ for $j=1,2,3,4$. The equality $t^{4}(v)=t^{4}(v^{\prime})$ implies that $u^{\prime}\in\{u,-u,\mathrm{i}u,-\mathrm{i}u\}$. In particular, if $u=0$, then $u^{\prime}=0$, and hence $v=v^{\prime}$. It remains to deal with the case $u\neq 0$. Moreover, $u=u^{\prime}$ if and only if $t(v)=t(v^{\prime})$ if and only if $t^{3}(v)=t^{3}(v^{\prime})$. Suppose for contradiction that $u\neq u^{\prime}$. Then $f_{j}(v)=f_{j}(v^{\prime})$ for $j=2,3,4$ and $(w_{1},w_{2})=(w^{\prime}_{1},w^{\prime}_{2})$ imply that
$0=r^{(1)}_{(1,-\mathrm{i})}(v)=r^{(2)}_{(1,-\mathrm{i})}(v)=r_{(1,\mathrm{i})}(v).$
From $0=r^{(1)}_{(1,-\mathrm{i})}(v)=r_{(1,\mathrm{i})}(v)$ we conclude $0=x_{1}(v)y_{2}(v)=x_{2}(v)y_{1}(v)$. Taking into account that $w_{1}\neq 0$ and $w_{2}\neq 0$, we conclude that $0=x_{1}(v)=y_{1}(v)$ or $0=x_{2}(v)=y_{2}(v)$. Then $r^{(2)}_{(1,-\mathrm{i})}(v)=0$ yields that both $x_{1}(v)y_{1}(v)$ and $x_{2}(v)y_{2}(v)$ are zero. We deduce $w_{1}=0$ or $w_{2}=0$, a contradiction. ∎
Proposition 8.4.
We have $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2}\oplus U_{(-1,1)})\leq 6$ and $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2}\oplus U_{(1,-1)})\leq 6$.
Proof.
First we deal with $V:=W_{1}\oplus W_{2}\oplus U_{(-1,1)}$. Let $v=(w_{1},w_{2},u)\in V$ and $v^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},u^{\prime})\in V$ such that $f(v)=f(v^{\prime})$ for all $f\in K[V]^{G}$ with $\deg(f)\leq 6$. We need to show that $v$ and $v^{\prime}$ have the same $G$-orbit. In the same way as in the proof of Proposition 8.3, we may assume that $w_{1}=w^{\prime}_{1}$, $w_{2}=w^{\prime}_{2}$ and none of $w_{1}$ or $w_{2}$ is zero. It follows that
$r_{(-1,1)}=x_{1}^{2}+x_{2}^{2},\quad r_{(-1,-1)}s_{(1,-1)}=(x_{1}^{2}-x_{2}^{2})y_{1}y_{2},\quad s_{(-1,1)}=y_{1}^{2}+y_{2}^{2}$
can not all vanish at $v$. Suppose that $f$ is one of the above polynomials, with $f(v)\neq 0$. Since $f$ is a relative invariant of weight $(-1,1)$ of degree at most $4$, we get that $ft\in K[V]^{G}$ has degree at most $5$. Moreover, the equality $(ft)(v)=(ft)(v^{\prime})$ implies $u=u^{\prime}$, hence $v=v^{\prime}$.
The proof for $V=W_{1}\oplus W_{2}\oplus U_{(1,-1)}$ is similar, the only difference is that we use the relative invariants
$s_{(1,-1)}=y_{1}y_{2},\quad s_{(-1,1)}r_{(-1,-1)}=(y_{1}^{2}+y_{2}^{2})(x_{1}^{2}-x_{2}^{2}),\quad s_{(-1,-1)}r_{(-1,1)}=(y_{1}^{2}-y_{2}^{2})(x_{1}^{2}+x_{2}^{2}).$
Proposition 8.5.
We have $\mu(G)\leq 2$, and hence $\kappa(G)\leq 3$.
Proof.
Let $S$ be an intersection independent set of subgroups of $G$ with $|S|=\mu(G)$. Clearly $G\notin S$. If $S$ contains a cyclic subgroup of prime power order, then $|S|\leq 2$ by Lemma 3.8. From now on assume that $S$ contains no cyclic subgroup of prime power order. All such proper subgroups of $G$ contain the subgroup $\langle a^{2},b^{2}\rangle$. If $\langle a^{2},b^{2}\rangle\in S$, then $|S|=1$. Otherwise $S$ consists of two order $8$ subgroups of $G$. In either case, $|S|\leq 2$. The inequality $\mu(G)\leq 2$ is proved, and hence $\kappa(G)\leq 1+\mu(G)\leq 3$ holds by (10). ∎
Theorem 8.6.
Assume that $K$ has an element of multiplicative order $8$. Then we have $\beta_{\mathrm{sep}}(\mathrm{C}_{4}\rtimes\mathrm{C}_{4})=6$.
Proof.
The factor group of $G$ modulo its central normal subgroup $\langle a^{2}b^{2}\rangle$ is isomorphic to the quaternion group $\mathrm{Dic}_{8}$ of order $8$. Therefore $\beta_{\mathrm{sep}}^{K}(G)\geq\beta_{\mathrm{sep}}^{K}(\mathrm{Dic}_{8})=6$ (in fact we have $\beta_{\mathrm{sep}}(G,W_{2})=6$, as we saw in Section 4).
To prove the reverse inequality, by Lemma 3.1 we may assume that $K$ is large enough so that Lemma 3.6 applies and we can reduce to the study of multiplicity-free representations. By Proposition 8.5 we have $\kappa(G)\leq 3$, and therefore by Lemma 3.7 it is sufficient to show that $\beta_{\mathrm{sep}}(V)\leq 6$ if $V$ is the direct sum of three pairwise non-isomorphic $G$-modules. Assume that $V$ is such a $G$-module. By Proposition 8.2, Proposition 8.1 we know that $\beta_{\mathrm{sep}}(G,V)\leq\beta(G,V)\leq 6$, unless $V=W_{1}\oplus W_{2}\oplus U_{\chi}$ for some non-trivial character $\chi\in\widehat{G}$. The cases when $\chi\in\{(1,\mathrm{i}),\ (-1,1),\ (1,-1)\}$ are covered by Proposition 8.3, Proposition 8.4. The remaining representations are of the form $\rho\circ\alpha$, where $\rho$ is one of the above three representations, and $\alpha$ is an automorphism of $G$. So the proof is complete by Lemma 3.11. ∎
8.2. The group $\mathrm{C}_{5}\rtimes\mathrm{C}_{4}$
@@
$f_{1}:=x_{1}x_{3}+x_{2}x_{4},\qquad f_{2}:=x_{1}x_{2}x_{3}x_{4},$
$f_{3}:=h_{1}+h_{2},\qquad f_{4}:=h_{1}h_{2},\qquad f_{5}:=(h_{1}-h_{2})h_{0},$
$f_{6}:=h_{3}+h_{4},\qquad f_{7}:=x_{1}^{5}+x_{2}^{5}+x_{3}^{5}+x_{4}^{5},$
$f_{8}:=x_{1}^{6}x_{3}+x_{1}x_{3}^{6}+x_{2}^{6}x_{4}+x_{2}x_{4}^{6}$
$f_{9}:=(h_{3}-h_{4})h_{0}.$
@@
Assume that $K$ has characteristic 00, and contains an element of multiplicative order $20$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{5}\rtimes\mathrm{C}_{4})=6$.
@@
The subgroup $\langle a,b^{2}\rangle$ of $G$ is isomorphic to $\mathrm{D}_{10}$, the dihedral group of order $10$, hence by equation (4) and Theorem 2.1 we have the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq 6$.
@@
(30) $f(v)=f(v^{\prime})\text{ holds for all homogeneous }f\in K[V]^{G}\text{ with }\deg(f)\leq 6.$
We have to show that $G\cdot v=G\cdot v^{\prime}$. By Proposition 8.8 we have $G\cdot w=G\cdot w^{\prime}$, so replacing $v^{\prime}$ by an appropriate element in its $G$-orbit we may assume that $w=w^{\prime}$. Moreover, $G/G^{\prime}=G/\langle a\rangle\cong\langle b\rangle\cong\mathrm{C}_{4}$, and $\mathsf{D}(\mathrm{C}_{4})=4$, so $G\cdot u=G\cdot u^{\prime}$ as well, implying that $u^{\prime}_{\mathrm{i}}\in\{\pm u_{\mathrm{i}},\pm\mathrm{i}u_{\mathrm{i}}\}$, $u^{\prime}_{-1}=\pm u_{-1}$, and $u^{\prime}_{1}=u_{1}$. So it is sufficient to deal with the case when $w\neq 0$ and $u_{\chi}\neq 0$ for some $\chi\in\widehat{G}\setminus\{1\}$.
@@
Case III.: $\mathrm{Stab}_{G}(w)=\{1_{G}\}$. Clearly it is sufficient to show that $u_{\chi}=u^{\prime}_{\chi}$ for all $\chi\in\widehat{G}\setminus\{1\}=\{-1,\pm\mathrm{i}\}$. First we deal with $\chi=-1$. Using the CoCalc platform [44] we obtained that a minimal homogeneous generating system of $K[W\oplus U_{-1}]^{G}$ consists of the generators $f_{1},\dots,f_{8}$ of $K[V]^{G}$ from Proposition 8.7, together with $t_{-1}^{2}$, $h_{0}t_{-1}$, $(h_{1}-h_{2})t_{-1}$, $(h_{3}-h_{4})t_{-1}$, $(x_{1}^{5}-x_{2}^{5}+x_{3}^{5}-x_{4}^{5})t_{-1}$. So all the generators involving $t_{-1}$ have degree at most $6$. This implies that $G\cdot(w,u_{-1})=G\cdot(w,u^{\prime}_{-1})$. Taking into account that the stabilizer of $w$ is trivial, this means that $u_{-1}=u^{\prime}_{-1}$.
@@
8.3. The group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$
In this section
$G=\mathrm{C}_{7}\rtimes\mathrm{C}_{3}=\langle a,b\mid a^{7}=b^{3}=1_{G},\quad bab^{-1}=a^{2}\rangle$
is the only non-abelian group of order $21$.
It is proved in [9, Theorem 4.1] that for any prime $p$ congruent to $1$ modulo $3$, the separating Noether number of the non-abelian semidirect product $\mathrm{C}_{p}\rtimes\mathrm{C}_{3}$ equals $p+1$, under the assumption that the base field $K$ contains an element of multiplicative order $3p$. Actually, [9] works over an algebraically closed base field, but the representation providing the exact lower bound for the separating Noether number is defined over any field containing an element of multiplicative order $3p$, and in order to prove that $p+1$ is also an upper bound, we may pass to the algebraic closure using Lemma 3.1.
As the special case $p=7$, we get the following:
Corollary 8.10.
Assume that $K$ contains an element of multiplicative order $21$. Then we have $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{7}\rtimes\mathrm{C}_{3})=8$.
8.4. The group $\mathrm{M}_{27}$

### v3

@@ 5. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
Case 2.a: $q_{yy}(w_{2})=0$ and $p^{-}_{yyy}(w_{2})\neq 0$. Then $w_{2}$ is a non-zero scalar multiple of $1,0]^{T}$ or $0,1]^{T}$, say $w_{2}$ is a non-zero scalar multiple of $1,0]^{T}$. If $w_{1}$ is a scalar multiple of $0,1]^{T}$, then none of $q_{xy}$, $q^{-}_{xy}$, $q_{xy}q^{-}_{xy}$, $q_{xy}^{2}$, $p^{-}_{yyy}$ vanishes on $v=(w_{1},w_{2})$, and these relative invariants on $V$ of degree at most $4$ exhaust all the five non-trivial weights, so we are done. Otherwise the first coordinate of $w_{1}$ is non-zero, and thus $p_{xxy}$, $p_{xyy}$, $p^{-}_{xxy}$, $p^{-}_{xyy}$ are relative invariants on $V$ of degree $3$ not vanishing on $v$ (just like $p^{-}_{yyy}$). So we have exhausted all the five non-trivial weights, and we are done.
Case 2.b: $p^{-}_{yyy}(w_{2})=0$ and $q_{yy}(w_{2})\neq 0$. Then $w_{2}$ is a non-zero scalar multiple of $1,1]^{T}$, $1,\omega]^{T}$, or $1,\omega^{2}]^{T}$. Moreover, $w_{1}$ is not a scalar multiple of $w_{2}$ (since otherwise its stabilizer would contain the stabilizer of $w_{2}$). Then $q_{yy}$, $q^{-}_{xy}$, $q_{yy}^{2}$, $q^{-}_{xy}q_{yy}$, $p^{-}_{xyy}$ are non-zero on $v=(w_{1},w_{2})$, and these relative invariants on $W_{1}\oplus W_{2}$ of degree at most $4$ exhaust all the five non-trivial weights.
Case 3: $V=W_{1}\oplus W_{3}$. The group $G$ has the automorphism $\alpha:a\mapsto a$, $b\mapsto b$, $c\mapsto c^{2}$. We have $\psi_{3}=\psi_{2}\circ\alpha$, so $\alpha$ interchanges $W_{2}$ and $W_{3}$, and permutes the $1$-dimensional representations. So Case 3 follows from Case 2 by Lemma 2.5 (i). ∎
@@
6. The group $\mathrm{C}_{5}\rtimes\mathrm{C}_{4}$
@@
$f_{1}:=x_{1}x_{3}+x_{2}x_{4},\qquad$ $f_{2}:=x_{1}x_{2}x_{3}x_{4},\qquad$ $f_{3}:=h_{1}+h_{2},$
$f_{4}:=h_{1}h_{2},\qquad$ $f_{5}:=(h_{1}-h_{2})h_{0},\qquad$ $f_{6}:=h_{3}+h_{4},$
$f_{7}:=x_{1}^{5}+x_{2}^{5}+x_{3}^{5}+x_{4}^{5},\qquad$ $f_{8}:=x_{1}^{6}x_{3}+x_{1}x_{3}^{6}+x_{2}^{6}x_{4}+x_{2}x_{4}^{6},\qquad$ $f_{9}:=(h_{3}-h_{4})h_{0}.$
@@
Assume that $K$ has characteristic 0, and contains an element of multiplicative order $20$. Then we have the equality $\beta_{\mathrm{sep}}^{K}(\mathrm{C}_{5}\rtimes\mathrm{C}_{4})=6$.
@@
The subgroup $\langle a,b^{2}\rangle$ of $G$ is isomorphic to $\mathrm{D}_{10}$, the dihedral group of order $10$, hence by inequality (4) and [15, Theorem 2.1] we have the inequality $\beta_{\mathrm{sep}}^{K}(G)\geq 6$.
@@
(20) f(v)=f(v) $f(v^{\prime})\text{ holds for all homogeneous }f\in K[V]^{G}\text{ with }\deg(f)\leq 6$.
[…]
@@
[…]
@@
7. The group $\mathrm{M}_{27}$